OpenStudy (anonymous):

Your weekly paycheck is 10 percent less than your coworker's. Both paychecks total 860 . Find the amount of each paycheck. The formula I got was a +[a + a(.10)] = 860 the answers i got was 450 and 410 but it says its not correct

OpenStudy (anonymous):

i tried 409.52 and 450.48 and its says thats incorret aslo

OpenStudy (anonymous):

y=x-0.1x x+y=860 substitute y into second equation x+x-0.1x=860 1.9x=860 x=860/1.9 452.631 plug in x value into one of the equation x+y=860 452.631+y=860 y=860-452.631 y=407.369

OpenStudy (jwt625):

x+0.9x=860,x=860/1.9=452.631

OpenStudy (anonymous):

so, Your weekly paycheck=407 and coworker's=453

OpenStudy (jwt625):

agree

OpenStudy (anonymous):

thank you so much thats correct